A) If an industry has exactly 20 firms with identical sales, the Herfindahl index must be:

1- less than 100.

2- greater than 100 but less than 200.

3- greater than 200 but less than 400.

4- greater than 400.

B) Which of the following is an argument in favor of judging competitiveness by performance rather than structure?

1- Judgment by performance requires that each action of a firm be analyzed on a case-by-case basis, which is very time-consuming and expensive.

2- Under the judgment by structure criterion, a firm may be breaking the law if it does what it's supposed to be doing: producing the best product it can at the lowest cost.

3- Courts know little economic theory.

4- Structure can be a predictor of future performance.
Answer #1

A) Option 4. Greater than 400

Explanation: All the 20 firms have the same market share, so, the market share of each firm = 100%/20 = 5%. Herfindahl index is the sum total of the square of the market share of each firm in the market. Here, the Herfindahl index is the sum total of the market share of each firm in the market. Here, the Herfindahl index = 20 * 52 = 20 * 25 = 500.